OpenStudy (♥ashley♥):

Solve for x. 3x^2 – 75 = 0 A. {25, –25} B. {25} C. {5} D. {5, –5}

jagr2713 (jagr2713):

Ok first we ad 75 on both sides right @♥Ashley♥

jagr2713 (jagr2713):

what will we have left

OpenStudy (♥ashley♥):

3x^2=0+75

jagr2713 (jagr2713):

no we add 5 on both sides s the 0 is gone right?

OpenStudy (♥ashley♥):

yes

jagr2713 (jagr2713):

so we are left with 3x2=75

jagr2713 (jagr2713):

What do we do next @♥Ashley♥

OpenStudy (♥ashley♥):

Divide 3x^2 on both sides

jagr2713 (jagr2713):

yes an what do u get

OpenStudy (kidrah69):

3x2−75=0 <---equation Add 75 to both sides 3x2=75 Divide both sides by 3 x2=75/3 Simplify 75/3 to 25 x2=25 Take the square root of both sides this way we get rid of x^2 x=±25 Simplify the square root of 25 to 5 x=±5

jagr2713 (jagr2713):

No we actually divide by 3 on both sides

jagr2713 (jagr2713):

@kidrah69 it better to guide them to the answer, not doing it for them

OpenStudy (kidrah69):

I did not give them the answer :) i showed my work

OpenStudy (kidrah69):

And they will better understand if it is split into steps :)

OpenStudy (♥ashley♥):

75/3=25

jagr2713 (jagr2713):

Do you understand

OpenStudy (♥ashley♥):

So the answer is 25 right

OpenStudy (kidrah69):

@♥Ashley♥ no whats the square root of 25

OpenStudy (♥ashley♥):

5

OpenStudy (kidrah69):

yeah its positive negative 5

OpenStudy (kidrah69):

since we square rooted it :) look at my work above

OpenStudy (♥ashley♥):

Thanks everyone! That was helpful!

OpenStudy (kidrah69):

=)